OpenStudy (anonymous):

Help would be much appreciated ^.^ -Find the absolute and local maximum and minimum values of f(t) = 2/t + 10 , 0 < t ≤ 4

myininaya (myininaya):

Well we know we need to find f'(t) is f(t)=2/t+10 or f(t)=2/(t+10)?

OpenStudy (anonymous):

the 1st one :)

myininaya (myininaya):

just so i can know to check your derivative

myininaya (myininaya):

ok cool

myininaya (myininaya):

\[f(t)=\frac{2}{t}+10 \\ f(t)=2t^{-1}+10 \\\] do you know how to find f'?

OpenStudy (anonymous):

Yes! It's \[f'(t)= \frac{ -2 }{ t^2 }\] am I right?

myininaya (myininaya):

yes and f' is not zero ever but f' dne at t=0 but we f(0) also doesn't exist at t=0 we actually have a vertical asymptote  so we don't have to worry about critical numbers but this just shows we also don't have any max of any type but w do have a min and it is an absolute min (you know since local min can't occur at endpoints)